OpenStudy (anonymous):

what is---- 5/3x - 1/y divided by 7x - 4/y?

OpenStudy (anonymous):

I find it's a good idea to first simplify complex fractions (though I'm not sure at all that this is what you have since your explanation of the problem isn't clear) \[{\frac{5}{3x} - \frac{1}{y} \over 7x - \frac{4}{y}}\cdot {3xy \over 3xy} = {5y - 3x \over 21x^2y - 12x}\]

OpenStudy (anonymous):

ty for ur answer